An athlete burns 9 calories per min on a stationary bicycle and 11 calories per min on a stair climber. In a 60 min workout the athlete burns 590 calories. How many minutes does the athlete spend on each type of exercise equipment?
Answer by ewatrrr(24785)

x the time on stair climber. Bicycle60-x
11x + 9(60-x) = 590
x = 50/2 = 25 min, time on stair climber. 35min on the Bicycle 60-25
...
Will Let You check the Answer
11(25) + 9(35)